OpenStudy (clara1223):

Give the values of A and B for the function f(x)={Ax-B when x<=1, -24x when 1= 5} to be continuous at both x = 1 and x = 5.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

this is not as hard as it may look replace \(x\) by \(1\) in both expressions \[Ax-B\] and \[-24x\]

OpenStudy (misty1212):

if \(x=1\)the first one is \(A-B\) and the second is \(-24\) so you have on equation \[A-B=-24\]

OpenStudy (misty1212):

then put in \(x=5\) in both \(-24x\) and \(Bx^2-A\)

OpenStudy (misty1212):

this time you get \[-120=25B-A\] as a second equation

OpenStudy (misty1212):

solve the system \[A-B=-24\\25B-A=-120\] to find \(A\) and \(B\)

OpenStudy (clara1223):

Thanks so much! I got that A= -30 and B=-6. I'm just going to plug them back in to make sure they work! ti\njhj

OpenStudy (misty1212):

i didn't do it, but i can check if you like

OpenStudy (misty1212):

yeah i get the same thing

OpenStudy (clara1223):

Don't worry about it, plugging it in will tell me if it's right or not.

OpenStudy (clara1223):

oh, thanks!

OpenStudy (misty1212):

\[f(x) = \left\{\begin{array}{rcc} Ax-B & \text{if} & x <1 \\ -24x& \text{if} & 1\leq x\leq5\\ Bx^2-A&\text{if} &x>5 \end{array} \right. \]

OpenStudy (misty1212):

just seeing if i could write it \[\color\magenta\heartsuit\]